Solve the differential equation = x dy / dx + 2y = x4

1
Expert's answer
2022-02-07T15:09:08-0500

The given equation can be written as

"\dfrac{dy}{dx} +\dfrac{2y}{x}= x^{3}" . This is of the form "\dfrac{dy}{dx}+Py=Q", whose solution is given by

"ye^{\int Pdx} = \int Qe^{\int Pdx} dx + c".


Here "P = \dfrac{2}{x}, Q= x^{3}" . Then, "e^{\int Pdx} = e^{\int \frac{2}{x}dx} = e^{2 \log x}= e^{\log x^{2}} = x^{2}." Thus the solution is

"\begin{aligned}
yx^{2} &= \int x^{3} \cdot x^{2} dx + c\\
yx^{2} &= \int x^{5} dx + c \\
yx^{2} &= \frac {x^{6}}{6} + c \\
\therefore y & = \frac {x^{4}}{6} + cx^{-2} \\
\end{aligned}"